OpenStudy (anonymous):

6 radians is the same as____ (round to the nearest hundredth of a degree) help please!

OpenStudy (mertsj):

\[6 radians \times \frac{180^{o}}{piradians}=\frac{6\times180^{o}}{3.14159}\]

OpenStudy (anonymous):

thanks!

OpenStudy (mertsj):

yw